How many gallons of 60% antifreeze needs to be mixed with 60 gallons of 20% antifreeze to make a mixture of 50% antifreeze? I am told to use the six step method and i am not too sure what that is.
Answer by jorel1380(3719)

n(.6)+60(.2)=(n+60)(.5)
.6n+12=.5n+30
.1n=18
n=18/.1=180 gallons of 60% antifreeze needed..